# A Game-Theoretic Framework for Distributed Load Balancing: Static and Dynamic Game Models

### v1

@@ 3 Main Theoretical Results for the Static Load Balancing Game
To that end, let us consider the potential function
@@
$=$ $\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{s_{ij}}{\mu_{j}}\right)^{2}+\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{\lambda_{i}^{2}a_{ij}^{2}}{\mu_{j}^{2}}+\frac{2\lambda_{i}a_{ij}s_{ij}}{\mu_{j}^{2}}\right)$
@@
where the last equality is obtained by the expression of the cost function (2) and the definition of $s_{ij}$. Using an identical argument, we have $\Phi(a^{\prime}_{i},a_{-i})=\sum_{j=1}^{m}\frac{\mu_{j}}{2}\big{(}\frac{s_{ij}}{\mu_{j}}\big{)}^{2}+D_{i}(a^{\prime}_{i},a_{-i})$. Finally, by subtracting this relation from the former one, we get the desired relation (6).
Since the static game is an exact potential game, it admits a pure NE. Moreover, any sequence of best-response updates, in which a player updates its strategy and strictly reduces its cost, will also reduce the potential function. As a result, the best-response dynamics will converge to a local minimum of the potential function, which must be a pure NE.
Remark 3.3.
@@
$\frac{\partial\Phi(a)}{\partial a_{ij}}=\frac{\partial D_{i}(a)}{\partial a_{ij}}=\lambda_{i}\big{(}\frac{\lambda_{i}a_{ij}+s_{ij}}{\mu_{j}}\big{)}\ \forall i,j.$
This in view of [43, Lemma 4.4] suggests that the static load balancing game is a potential game.
Remark 3.4.
One can show that the potential function $\Phi(a)$ in the proof of Theorem 1 is strongly convex with respect to each player’s action separately. However, in general, $\Phi(a)$ may not be jointly convex with respect to the entire action profile $a$. Otherwise, we would have a unique pure NE.
@@
$\frac{\partial^{2}D_{i}(a)}{\partial a_{ij}\partial a_{i\ell}}=\begin{cases}\frac{\lambda_{i}^{2}}{\mu_{j}}\ &\mbox{if}\ \ell=j\\ 0\ &\mbox{if}\ \ell\neq j.\end{cases}$
@@
$\hat{\mu}_{ic_{i}}\left(\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ij}}\right)\leq\sum_{k=1}^{c_{i}-1}\mu_{k}.$
Proof 3.6.
@@
where $\alpha$ and $\eta=(\eta_{1},\ldots,\eta_{m})$ are Lagrange multiplies corresponding to the constraints in (8). Since (8) is a convex program, using KKT optimality conditions, we know that $a_{i}$ is the optimal solution (for a fixed $a_{-i}$) if and only if:
$\frac{\partial L_{i}}{\partial a_{ij}}=0\ \forall j,\quad\frac{\partial L_{i}}{\partial\alpha}=0,$
@@
Now, a key observation is that a server with a higher available processing rate should naturally have a higher fraction of jobs assigned to it. Under the assumption that the servers are sorted by their available processing rates, the load fractions on the servers also should follow the same ordering, i.e.,
$a_{i1}\geq a_{i2}\geq\cdots\geq a_{im}.$
Consequently, there may be cases where slower servers are assigned no jobs. This implies the existence of an index $c_{i}\in[n]$ such that
$a_{ij}=0\quad\text{for }i=c_{i},\ldots,n.$
@@
where $c_{i}\in[n]$ is the smallest index satisfying
@@
Theorem 3.7.
Suppose in the static load balancing game each player is initially distributing a non-zero portion of its job on each server, i.e., $a_{ij}^{0}\neq 0,\ \forall i,j$. Then, the action profile obtained when players update their actions exactly once (in any order) using the best-response rule in (11) is a pure NE. In other words, the best-response dynamics converge to a pure NE in $n$ iterations.
Proof 3.8.
@@
4 Main Theoretical Results for the Dynamic Load Balancing Game
@@
Upon arrival of a new job at time $t$ to some player $i$ with job length $\lambda_{i}$:
  • Player $i$ observes the state $s^{t}$ at time $t$, computes the available processing rates $\hat{\mu}^{t}_{j}=\frac{\mu_{j}}{s_{j}^{t}},j\in[m]$, and sorts them in an increasing order, where without loss of generality, we may assume $\hat{\mu}^{t}_{1}\geq\cdots\geq\hat{\mu}^{t}_{m}.$
  • Player $i$ computes its best response at time $t$ by solving the convex program $\min\{D_{i}(a_{i},s^{t}):a_{i}\in A_{i}\}$, whose solution is given by
    $a^{t}_{ij}=\begin{cases}\frac{\mu_{j}}{\lambda_{i}}\left(\frac{\lambda_{i}+\sum_{k=1}^{c-1}s^{t}_{k}}{\sum_{k=1}^{c-1}\mu_{k}}-\frac{s^{t}_{j}}{\mu_{j}}\right)&1\leq j<c,\\[10.0pt] 0&c\leq j\leq m,\end{cases}$
    where $c$ is the smallest index satisfying:
    $\hat{\mu}^{t}_{c}\left(\lambda_{i}+\sum_{k=1}^{c-1}s_{k}\right)\leq\sum_{k=1}^{c-1}\mu_{k}.$
  • Player $i$ schedules $\lambda_{i}a^{t}_{ij}$ portion of its job on server $j\in[m]$, and the game proceeds to the next step $t+1$.
@@
Therefore, according to (4), the states of the servers evolve as
@@
To show that the nested sequence (20) must grow, we first note that if a server $j$ becomes empty at some time $t$, i.e., $s_{j}^{t}=0$, then $\hat{\mu}_{j}^{t}=\infty$. Thus, server $j$ must receive a job at the next time step $t+1$, and so by the nested property (20), it must receive a job in all subsequent iterations. Since $c\notin S^{t}$, it means that in all iterations $1,2,\ldots,t$, server $c$ did not receive a new job and was only processing its initial load $s^{0}_{c}$ at rate $\mu_{c}$. Otherwise, server $c$ must have previously received a job and thus would be in the set $S^{t}$ by the nested property (20). Therefore, the load of server $c$ at time $t$ would be $s_{c}^{0}-t\mu_{c}$. As a result, server $c$ does not belong to $S^{t}$ for at most $t\leq\lceil s^{0}_{c}/\mu_{c}\rceil$ iterations, after which its load becomes zero and according to the above discussion, it will receive a job in the next iteration. By repeating this argument for any choice of server $c$, one can see that after at most $t^{\prime}=\max_{j\in[m]}\lceil s^{0}_{j}/\mu_{j}\rceil$ iterations, we must have $S^{t^{\prime}}=[m]$.
We are now ready to state the main result of this section.
Theorem 4.11.
@@
$t^{\prime\prime}=O\Big{(}\max_{j}\lceil\frac{s^{0}_{j}}{\mu_{j}}\rceil\frac{\sum_{j}\mu_{j}+\sum_{j}s_{j}^{0}}{\sum_{j}\mu_{j}-\lambda_{\max}}\Big{)}$
@@
Proof 4.12.
From Lemma 4.9, we know that once a player begins sending jobs to all servers, all subsequent players will continue to do so for the remainder of the time. Consider an arbitrary time $t\geq t^{\prime}$, and assume that at time $t$ player $i$ receives a job and plays its best-response according to Algorithm 1. Since all servers receive jobs at time $t\geq t^{\prime}$, we have
@@
  • If $C^{t}>1$, then $\sum_{j=1}^{m}s_{j}^{t+1}\neq 0$. As in (16) where $c$ is replaced by $m$, we have
@@
$t^{\prime\prime}$ :-$\coloneq t^{\prime}+\frac{\sum_{j=1}^{m}s^{t^{\prime}}_{j}}{\sum_{j=1}^{m}\mu_{j}-\lambda_{\max}}:$
@@
Finally, we note that since $\sum_{j=1}^{m}s^{t}_{j}=0$ for $t\geq t^{\prime\prime}$, the dynamic load balancing game reduces to the static load balancing game with zero initial loads, i.e., $s^{0}_{j}=0\ \forall j\in[m]$. Since by Theorem 1, the best-response strategy in the static game converges to a pure NE, we obtain that $\lim_{t\to\infty}a^{t}_{ij}=\frac{\mu_{j}}{\mu_{1}+\dots+\mu_{m}}$ for any $i\in[n]$ and $j\in[m]$, which is precisely the pure NE of the static game with zero initial loads.
@@
Parameters for Setting 1: $\mu=[0.9,0.8,0.4,0.01]$, $\lambda=[0.5,0.5,0.3,0.7]$, $s^{0}=[10,10,1,0.5]$.
(d) One player distributes a job at a time.
(e) All players distribute jobs simultaneously.
(f) Sequentially playing best-response dynamics.
Parameters for Setting 2: $\mu=[1.4,1.2,1,0.5]$, $\lambda=[0.5,0.5,0.3,0.7,0.9,0.1,0.6,0.2]$, $s^{0}=[10,10,1,10]$.
(g) One player distributes a job at a time.
(h) All players distribute jobs simultaneously.
(i) Sequentially playing best-response dynamics.
Parameters for Setting 3: $\mu=[2,2]$, $\lambda=[0.5,0.5,0.3,0.7,0.9,0.1,0.6,0.2]$, $s^{0}=[10,50]$.
Figure 2: Comparison of different job distribution strategies across three settings. Each column corresponds to a specific strategy, and each row corresponds to a specific setting. Parameters for each row are included in their respective captions.
(a) $\mu=[1,1.5,2,0.6]$, $\lambda=[5,5,5,5]$, $s^{0}=[10,10,10,10]$
(b) $\mu=[0.3,0.8,0.2,0.01]$, $\lambda=[1,0.8,0.9,0.7]$, $s^{0}=[10,10,1,0.5]$
@@
5 Numerical Experiments
In this section, we present numerical results to evaluate the performance of the best-response dynamics under various settings of the static game, dynamic game, sequential update rule, and simultaneous update rule. In our simulations, we consider three different choice of parameters for the initial state, players’ job lengths, and servers’ service rates, where the parameters for each setting are provided in the captions of the corresponding figures. For instance, in the first row of Fig. 2, we consider a settings with four servers with initial loads $s^{0}_{1}=10,s_{2}^{0}=10,s_{3}^{0}=1,s_{4}^{0}=0.5$ and service rates $\mu_{1}=0.9,\mu_{2}=0.8,\mu_{3}=0.4,\mu_{4}=0.01$, and a set of four players with job lengths $\lambda_{1}=0.5,\lambda_{2}=0.5,\lambda_{3}=0.3,\lambda_{4}=0.7$.
First, we simulate the performance of Algorithm 1 for the dynamic load balancing game, where at each time step, a single player is randomly selected to distribute its job across the servers. The results of this experiment are illustrated in Figs. 2(a), 2(d), 2(g), 3(a), and 3(b). As can be seen, after sufficient time has elapsed, all states converge to zero. Once the load on all servers becomes zero, it remains at zero, indicating a stable NE. As illustrated in these figures, the convergence of the dynamics to the zero state appears to be fast and scales polynomially with respect to the game parameters, which also justifies the theoretical performance guarantee provided by Theorem 4.11.
Next, we evaluate an alternative version of Algorithm 1 for the dynamic load balancing game, in which, at each time step, all players simultaneously distribute their jobs. In this case, each player adopts the best-response strategy, considering both the current state of the system and the strategies of other players. The results of this experiment are depicted in Figs. 2(b), 2(e), and 2(h). Interestingly, the results show that the simultaneous update of the players does not hurt the overall performance of the system and the state of the game eventually converges to zero, albeit at a slower convergence rate.
Finally, in Figs. 2(c), 2(f), and 2(i), we evaluate the performance of the sequential best-response dynamics in the static load balancing game, where $a^{0}_{ij}=\frac{1}{m}$ for all $i\in[n],j\in[m]$. As can be seen, the action profiles converge to a pure Nash equilibrium (NE) after exactly $n$ iterations (one update by each player), which confirms the theoretical convergence rate bound provided in Theorem 3.7.

### v2

@@ 3 Theoretical Results for the Static Game
In particular, the game admits a pure NE, and the best-response dynamics converge to a pure NE.
{pf}
Let us consider the potential function
@@
[…]
Remark 3.
@@
$\frac{\partial\Phi(a)}{\partial a_{ij}}=\frac{\partial D_{i}(a)}{\partial a_{ij}}=\lambda_{i}\left(\frac{\lambda_{i}a_{ij}+s_{ij}}{\mu_{j}}\right)\ \forall i,j.$
This in view of [49, Lemma 4.4] suggests that the static load balancing game is an exact potential game.
Remark 4.
One can show that the potential function $\Phi(a)$ in the proof of Theorem 2 is strongly convex with respect to each player’s action separately. However, in general, $\Phi(a)$ may not be jointly convex with respect to the entire action profile $a$.
@@
$\frac{\partial^{2}D_{i}(a)}{\partial a_{ij}\partial a_{i\ell}}=\begin{cases}\frac{\lambda_{i}^{2}}{\mu_{j}}\ &\mbox{if}\ \ell=j,\\ 0\ &\mbox{if}\ \ell\neq j.\end{cases}$
@@
$\hat{\mu}_{ic_{i}}\left(\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ik}}\right)\leq\sum_{k=1}^{c_{i}-1}\mu_{k}.$ (12)
{pf}
@@
where $\alpha$ and $\eta=(\eta_{1},\ldots,\eta_{m})$ are Lagrange multipliers corresponding to the constraints in (8). Since (8) is a convex program, using KKT optimality conditions, we know that $a_{i}$ is the optimal solution (for a fixed $a_{-i}$) if and only if:
$\frac{\partial L_{i}}{\partial a_{ij}}=0\quad\forall j,\quad\frac{\partial L_{i}}{\partial\alpha}=0,$
@@
Now, a key observation is that a server with a higher available processing rate should naturally have a higher fraction of jobs assigned to it. Under the assumption that the servers are sorted by their available processing rates, the load fractions on the servers also should follow the same ordering, i.e., $a_{i1}\geq a_{i2}\geq\cdots\geq a_{im}$. Consequently, there may be cases where slower servers are assigned no jobs. This implies the existence of an index $c_{i}\in[m]$ such that
$a_{ij}=0\quad\text{for }j=c_{i},\ldots,m.$
@@
where $c_{i}\in[m]$ is the smallest index satisfying
@@
[…]
Remark 6.
In practice, players may not have accurate estimates of the game parameters. To assess the robustness of best-response dynamics under observation noise, suppose the service rates are observed with noise: $\tilde{\mu}_{k}=\mu_{k}+\epsilon_{k}$, where $|\epsilon_{k}|\leq 1$. In this setting, one can show that the resulting best-response deviations are bounded in terms of the servers’ service rates, provided the noise is small enough not to alter $c_{i}$ as defined in Eq. (12).
@@
Theorem 7.
Suppose in the static load balancing game each player is initially distributing a non-zero portion of its job on each server, i.e., $a_{ij}^{0}\neq 0,\ \forall i,j$. Then, the action profile obtained when players update their actions exactly once (in any order) using the best-response rule in Lemma 5 is a pure NE. In other words, the best-response dynamics converge to a pure NE in $n$ iterations.
{pf}
Please see Appendix A.1 for the proof. We now proceed to bound the Price of Anarchy (PoA) of the static load balancing game, which is defined as the ratio between the worst-case social cost at a pure NE and the optimal social cost:
$\text{PoA}=\frac{\max_{a\in\text{NE}}\sum_{i=1}^{n}D_{i}(a)}{\min_{a}\sum_{i=1}^{n}D_{i}(a)}.$
The PoA quantifies the inefficiency of Nash equilibria in games due to the selfish behavior of the players, compared to a globally optimal solution [50].
Theorem 8.
The price of anarchy of the static load balancing game is bounded by
$\text{PoA}\leq 1+2\left(\max_{j\in[m]}\frac{s_{j}^{0}}{\mu_{j}}+\frac{\sum_{i=1}^{n}\lambda_{i}}{\mu_{\mathrm{min}}}\right)\left(\frac{\sum_{j=1}^{m}\mu_{j}}{\sum_{i=1}^{n}\lambda_{i}}\right).$
{pf}
@@
Remark 9.
In fact, the bound obtained in Theorem 8 can be further improved for special cases, such as when the initial loads are all zero, i.e., $s_{j}^{0}=0$ for all $j\in[m]$. In this case, as a direct corollary of the proof of Theorem 8, one can show that $\text{PoA}\leq 3$. We refer to Corollary A.3 in Appendix A.2 for a formal proof of this result.
4 Theoretical Results for the Dynamic Game
@@
Arrival of a new job at time $t$ to player $i$ with length $\lambda_{i}$:
  • Player $i$ observes state $s^{t}$ and computes available rates $\hat{\mu}^{t}_{j}=\frac{\mu_{j}}{s_{j}^{t}},\,j\in[m]$.
  • Sort servers so that $\hat{\mu}^{t}_{1}\geq\hat{\mu}^{t}_{2}\geq\cdots\geq\hat{\mu}^{t}_{m}$.
  • Compute best response $a_{i}^{t}$ by solving $\min\{D_{i}(a_{i},s^{t}):a_{i}\in A_{i}\}$, using Eqs. (11) and (12).
  • Schedule portions $\lambda_{i}a^{t}_{ij}$ of the job on server $j\in[m]$.
  • Proceed to next time step $t+1$.
@@
Therefore, according to Eq. (4), the states of the servers evolve as
@@
[…]
Theorem 11.
@@
$t^{\prime\prime}=O\left(\max_{j}\lceil\frac{s^{0}_{j}}{\mu_{j}}\rceil\frac{\sum_{j}\mu_{j}+\sum_{j}s_{j}^{0}}{\sum_{j}\mu_{j}-\lambda_{\max}}\right)$
@@
{pf}
From Lemma 10, we know that once a player begins sending jobs to all servers, all subsequent players will continue to do so for the remainder of the time. Consider an arbitrary time $t\geq t^{\prime}$, and assume that at time $t$ player $i$ receives a job and plays its best-response according to Algorithm 1. Since all servers receive jobs at time $t\geq t^{\prime}$, we have
@@
  • If $C^{t}>1$, then $\sum_{j=1}^{m}s_{j}^{t+1}\neq 0$. As in Eq. (17) where $c$ is replaced by $m$, we have
@@
$t^{\prime\prime}$ $\mathrel{:=}t^{\prime}+\frac{\sum_{j=1}^{m}s^{t^{\prime}}_{j}}{\sum_{j=1}^{m}\mu_{j}-\lambda_{\max}}$
@@
[…]
Remark 12.
Note that the bound on $t^{\prime\prime}$ given in Theorem 11 is a worst-case estimate. In practice, for many parameter settings, the system may converge significantly faster than the theoretical bound suggests (see Table 2). Appendix A.3 provides an alternative bound that can be tighter for certain parameter ranges.
Remark 13.
Theorem 11 shows that under the best-response policies (Algorithm 1), server loads vanish in polynomial time, and each player’s overall cost converges to the unique minimizer of their zero-state instantaneous cost (Eq. (3)). Since the overall cost is defined as the time-average of instantaneous costs (Eq. (5)), the cost for player $i$ converges to $\lambda_{i}^{2}/(2\sum_{j=1}^{m}\mu_{j})$. Thus, no player has an incentive to deviate, and the best-response policies form a NE, yielding a fully load-balanced system.
5 Numerical Experiments
@@
Figure 4: Comparison of different job distribution strategies for Setting 1, with parameters: $\mu=[1.4,1.4,1.2,0.5,0.4,0.3,0.2,0.1],\lambda=[1.5,0.5,0.3,0.7,\\ 0.9,0.1,0.6,0.2],s^{0}=[10,10,1,10,20,20,10,1]$.
(a) Average server load under sequential and simultaneous updates of the dynamic game.
(b) Average player cost in the static game.
Figure 5: Parameters for Setting 2: 500 players, 200 servers. Service rates $\sim\mathcal{U}[3,4]$, job lengths $\sim\mathcal{U}[2,3]$. Initial server states are sampled uniformly from $[10,20]$.
In this section, we present numerical results to evaluate the performance of the best-response dynamics under various settings of the static game, dynamic game, sequential update rule, and simultaneous update rule. In our simulations, we consider different choice of parameters for the initial state, players’ job lengths, and servers’ service rates, where the parameters for each setting are provided in the captions of the corresponding figures. For instance, in Fig. 4, we illustrate Setting 1, which involves eight servers with initial loads $[10,10,1,10,20,20,10,1]$ and service rates $[1.4,1.4,1.2,0.5,0.4,0.3,0.2,0.1]$, and a set of eight players with job lengths $[1.5,0.5,0.3,0.7,0.9,0.1,0.6,0.2]$.
First, we simulate the performance of Algorithm 1 for the dynamic load balancing game, where at each time step, a single player is randomly selected to distribute its job across the servers. The results of this experiment is illustrated in Fig. 4(a). As can be seen, after sufficient time has elapsed, all states converge to zero. Once the load on all servers becomes zero, it remains at zero, indicating a stable NE. As illustrated in this figure, the convergence of the dynamics to the zero state appears to be fast and scales polynomially with respect to the game parameters, which also justifies the theoretical performance guarantee provided by Theorem 11.
Next, we evaluate an alternative version of Algorithm 1 for the dynamic load balancing game, in which, at each time step, all players simultaneously distribute their jobs. In this case, each player adopts the best-response strategy, considering both the current state of the system and the strategies of other players. The result of this experiment is depicted in Fig. 4(b). Interestingly, the results show that the simultaneous update of the players does not hurt the overall performance of the system and the state of the game eventually converges to zero, albeit at a slower convergence rate.
Finally, in Fig. 4(c), we evaluate the performance of the sequential best-response dynamics in the static load balancing game, where $a^{0}_{ij}=\frac{1}{m}$ for all $i\in[n],j\in[m]$. As can be seen, the action profiles converge to a pure Nash equilibrium (NE) after exactly $n$ iterations (one update by each player), which confirms the theoretical convergence rate bound provided in Theorem 7.
To demonstrate the effectiveness of our framework in real-world scenarios, we evaluate an additional setting with significantly larger numbers of players and servers. Specifically, Setting 2 includes 500 players and 200 servers. In this case, each player’s job lengths are drawn independently from the uniform distribution on $[2,3]$, while the service rates of the servers are drawn uniformly from $[3,4]$. Additionally, the initial state of each server is initialized by sampling uniformly from the interval $[10,20]$. For this setting, we plot the average load on servers over time and compare the results for both sequential and simultaneous updates in the dynamic game. The results are depicted in Fig. 5(a). Furthermore, we plot the average player cost over time for the sequential update in the static load balancing game, which is shown in Fig. 5(b).
It is also worth noting that certain parameter settings are feasible under the sequential update of the dynamic game but not under the simultaneous update. For instance, consider a system with service rates $[3,2,1.5,1.4,1,0.5,0.2,0.1]$ and job lengths $[2.5,3,4,6,5,5,5,5]$. This setting is feasible for sequential updates because, for each player $i$, we have $\lambda_{i}<\sum_{j=1}^{8}\mu_{j}=9.7$. However, it becomes infeasible under simultaneous updates since the total arrival rate exceeds the total service rate: $\sum_{i=1}^{8}\lambda_{i}=35.5>\sum_{j=1}^{8}\mu_{j}=9.7$. Figure 6 illustrates the performance of Algorithm 1 for two such cases.
(a) $\mu=[3,2,1.5,1.4,1,0.5,0.2,0.1]$, $\lambda=[2.5,3,4,6,5,5,5,5]$, $s^{0}=[10,10,10,10,5,4,4,1]$
(b) $\mu=[1.5,1.2,1.0,0.8,0.3,0.2,0.1,0.5\\ ,0.4,0.01]$, $\lambda=[4,0.8,0.9,0.7,0.6,0.5,0.4,0.3,3]$, $s^{0}=[10,10,1,5,5,4,4,0.1,0.1,0.5]$
@@
Next, we further extend our numerical results by systematically varying the number of players ($n$) and servers ($m$), and analyzing how the convergence time of Algorithm 1 scales with system size. In this experiment, both $n$ and $m$ are varied over the set $\{20,40,60,80\}$. The service rates $\mu_{j}$ are sampled independently and uniformly from the interval $[1,2]$, while job lengths are drawn uniformly from $[0,1]$. The initial queue states are also initialized by sampling uniformly from $[10,20]$. The results of this experiment are shown in Fig. 7.
Figure 7: Convergence time for varying numbers of players and servers.
Finally in Table 2 we compare the iteration counts of sequential and simultaneous update strategies across various settings in the dynamic game. Settings 1 and 2 are defined in Figs. 4 and 5 respectively. Settings 3-7 are defined as follows:
  • Setting 3: 100 players, 50 servers. Service rates $\sim\mathcal{U}[1,2]$, job lengths $\sim\mathcal{U}[0,1]$, initial states $\sim\mathcal{U}[10,20]$.
  • Setting 4: 200 players, 100 servers. Service rates $\sim\mathcal{U}[2,3]$, job lengths $\sim\mathcal{U}[1,2]$, initial states $\sim\mathcal{U}[10,20]$.
  • Setting 5: $\mu=[0.9,0.8,0.4,0.01]$, $\lambda=[0.5,0.5,0.3,0.7]$, $s^{0}=[10,10,1,0.5]$.
  • Setting 6: $\mu=[1.4,1.2,1,0.5]$, $\lambda=[0.5,0.5,0.3,0.7,0.9,\\ 0.1,0.6,0.2]$, $s^{0}=[10,10,1,10]$.
  • Setting 7: $\mu=[2,2]$, $\lambda=[0.5,0.5,0.3,0.7,0.9,0.1,0.6,\\ 0.2]$, $s^{0}=[10,50]$.
Setting Iter. count (simult.) Iter. count (seq.)
1 110 66
2 6 6
3 32 19
4 9 8
5 176 45
6 90 19
7 280 24
Table 2: Iteration counts for simultaneous vs. sequential update strategies under various settings.